OpenStudy (anonymous):

the dim(u)=6 is understandable but dim(S)=6.....I don't get it. Can some body tell me the basis for symmetric 3X3 matrices

OpenStudy (anonymous):

dimension just means the number of vectors in the basis. one basis i can think of is \[\left[\begin{matrix}1 & 0 & 0 \\0 & 0 & 0 \\0 & 0 & 0 \end{matrix}\right],\left[\begin{matrix}0 & 0 & 0 \\0 & 1 & 0 \\0 & 0 & 0 \end{matrix}\right],\left[\begin{matrix}0 & 0 & 0 \\0 & 0 & 0 \\0 & 0 & 1 \end{matrix}\right],\left[\begin{matrix}0 & 1 & 0 \\1 & 0 & 0 \\0 & 0 & 0 \end{matrix}\right],\left[\begin{matrix}0 & 0 & 1 \\0 & 1 & 0 \\1 & 0 & 0 \end{matrix}\right],\left[\begin{matrix}0 & 0 & 0 \\0 & 0 & 1 \\0 & 1 & 0 \end{matrix}\right]\]